# Melbourne, VIC 3010, Australia

### v2

@@ 7.1 Minkowski vacuum and the 4d prepotential
We briefly highlight key aspects of 4d ${\cal N}=2$ supergravity to clarify its connection to the equivariant topological strings framework. The 4d action, whether in superspace or component form, is governed by the prepotential, a holomorphic function of the complex scalars $X^{I}$ [^28] In off-shell supergravity, the scalars $X^{I}$ are the lowest components of the vector multiplets and the vector superfield in superspace. On-shell, they overparametrize the physical scalars but are still used to define the theory via the prepotential. Supersymmetry ensures that the prepotential fully determines the Lagrangian, including higher-derivative corrections. It must also be a homogeneous function of a specific degree, related to the number of derivatives.
@@
where $C_{abc}$ are the triple intersection numbers of the CY three-fold, cf. (15) - (16), with $a,b,c=1,\dots,h^{1,1}(\mathrm{CY}_{3})$, and we assume an implicit summation over repeated up-down indices. The additional scalar $X^{0}$ (we use the convention $I=\{0,a\}$) in the denominator comes precisely from the circle reduction from 5d to 4d, and gives us a strong hint about the relation to equivariant topological strings on ×$\mathbb{C}\times\mathrm{CY}_{3}$. First, we can evaluate the equivariant volume of $X$, which clearly has a product structure
@@
We move on to the second term in the derivative expansion of the prepotential, which supersymmetry constrains to be of degree 00 in the scalars $X^{I}$ and linear in an additional composite scalar $\hat{A}$ [^30] For more details on the superconformal formalism that suitably accounts for higher-derivative corrections, see e.g. Lauria:2020rhc . The scalar $\hat{A}$ in this case determines the coupling to the so-called Weyl2 four-derivative invariant, Bergshoeff:1980is .
$F^{4\partial}(X;\hat{A})=\frac{1}{24}\,\frac{b_{a}X^{a}}{X^{0}}\,\hat{A}\ ,$ (233)
@@ 7.2 Black holes and the OSV conjecture
(×$\begin{array}[]{ccc}(\mathbb{C}\times\mathrm{CY}_{3})&\to&Y\\ &&\downarrow\\ &&\mathbb{P}^{1}\end{array}$ (239)
@@
Our equivariant approach proves useful here, as it explains why a formula like (240) should hold for the topological string from the 11d perspective. We are also able to partially prove this result in the toric directions.
@@
where we denoted the equivariant parameter on the two-sphere as $\bar{\nu}=a\epsilon_{-}-b\epsilon_{+}$ (with $a=b=1$ on the sphere), as we reserve the label 00 for the parameters coming from the $\mathbb{C}$ direction in the fiber as in the previous subsection. The equivariant volume of ×$\mathbb{C}\times\mathrm{CY}_{3}$ was already evaluated there, (230), and just as in Section 5 we denoted appropriate shift of the relevant parameters on the two poles with the $\pm$ subscript. On the complex plane we explicitly derived the shifts in $\lambda^{0}$ and $\epsilon_{0}$ in Subsection 5.1:
@@
The above conjecture for the shifts $t^{a}_{\pm}$ can be viewed as a reformulation of the OSV conjecture. It seems plausible that the shifts $t^{a}_{\pm}$ here come from the complexification of the real Kähler parameters by the reduction of the 11d three-form field, $C_{3}$. On the other hand, we have rigorously derived the shift along the toric direction ϵ0±superscriptsubscriptitalic-$\epsilon_{0}^{\pm}$, finding perfect agreement with higher-derivative supergravity.
@@ 8 Discussion and outlook
We finish with a discussion that relates directly to the topological string partition function discussed here. We give a list of topics more closely related to supergravity and holography in Cassia:2024 .
@@
$F_{X}(\lambda,\epsilon;g_{s},\mathfrak{b})=-\sum_{\mathfrak{g}=0}^{\infty}\sum_{n=0}^{\mathfrak{g}}g_{s}^{2(\mathfrak{g}-1)}\,(\mathfrak{b}+\mathfrak{b}^{-1})^{2n}F_{\mathfrak{g},n}(\lambda,\epsilon)\ ,$ (253)
@@
An apparent caveat of the refinement is the lack of a clear geometric origin for the additional parameter at the level of the CY manifold $X$. This particularly affects the constant maps terms, whose refinement lacks a general definition. We hope to come back to this point in the future, Cassia:2024 .
@@
where the power of $\omega$ gives the homogeneity degree in the moduli $\lambda^{i}$. Each of these three candidates correspond to one of the terms in (56). This is a choice that appears specifically relevant in the geometric description of M-theory, but this is no longer true for other brane models, see in particular (1.7) of Colombo:2023fhu .
@@
where $N_{\rm M2}$ and $N_{\rm M5}$ correspond to the exact brane charge of the corresponding type. We should however note that the relation between the transverse space $L$ and the modelled $X$ might generically become more complicated, see some examples in Martelli:2023oqk .
Similarly, it is easy to imagine a different equivariant topological string definition, relating to double intersections numbers, where the constant maps terms will be related to the equivariant upgrade of
@@
which would be used in the description of D3-branes and other type IIB objects,
@@
where the leading term in $F^{(2)}_{X}$ is proportional to $\mathbb{V}^{(2)}_{X}(\lambda,\epsilon)$, cf. (45), instead of $\mathbb{V}^{(3)}_{X}(\lambda,\epsilon)$.
Likewise, we can go to an arbitrary degree $m$,
@@
with each higher degree allowing a larger number of different invariants made up from $m$ derivatives acting on the equivariant volume $\mathbb{V}_{X}$ (we skipped $c_{1}(TX)$ as it vanishes on CY manifolds non-equivariantly), and first term proportional to $\mathbb{V}_{X}^{(m)}$. Such an example of higher degree $({\rm deg}=5)$ is a candidate for the description of massive IIA objects, see again (1.7) of Colombo:2023fhu ,
@@
Note furthermore that each of the terms of a given degree might allow for additional refinement and come in a different linear combination, see the discussion of periods in Cassia:2022lfj . We give additional holographic comments on these observations in Cassia:2024 .
@@ Acknowledgements
We wish to thank Thorsten Hertl for illuminating discussions, and Dario Martelli and Alberto Zaffaroni for useful comments on the draft. We gratefully acknowledge support during the MATRIX Program “New Deformations of Quantum Field and Gravity Theories,” (Creswick AU, 22 Jan – 2 Feb 2024), where this project was initiated. The study of KH is financed by the European Union- NextGenerationEU, through the National Recovery and Resilience Plan of the Republic of Bulgaria, project No BG-RRP-2.004-0008-C01.
Appendix A $\mathbb{P}^{k}$ and $\mathbb{C}^{m}\times\mathbb{P}^{k}$ fibrations
@@
$Q=\begin{pmatrix}1&1&0&-k\\ 0&0&1&1\end{pmatrix}\ ,$ (260)
@@
$Q=\begin{pmatrix}a&b&0&-k\\ 0&0&1&1\end{pmatrix}\ ,$ (268)
@@
$Q=\begin{pmatrix}a&b&0&\dots&0&-k\\ 0&0&1&\dots&1&1\end{pmatrix}\ ,$ (270)
@@
×$\mathbb{C}\times\mathbb{P}^{1}$ over $\mathbb{P}^{1}$
Let us consider the fiber ×$\mathbb{C}\times\mathbb{P}^{1}$ over a $\mathbb{P}^{1}$ base, giving us an interesting example of a product space fiber with a compact and a non-compact factor. This is the closest example to the situation discussed in the main text in section 7.2 that features a non-toric compact factor in the fiber.
@@
$Q=\begin{pmatrix}1&1&0&-n&-p\\ 0&0&1&1&0\end{pmatrix}\ ,$ (272)
@@
It is particularly interesting to observe what happens in the non-equivariant limits in this case. Setting to zero the equivariant paramters $\epsilon_{3},\epsilon_{4}$ associated to the compact directions on the fiber, we find
$\mathrm{vol}(t,\epsilon,\epsilon_{1,2},\epsilon_{3,4}=0)=\frac{1}{\epsilon_{2}-\epsilon_{1}}\,\left(\frac{e^{-t^{1}\epsilon_{1}}(1-e^{-n\,t^{2}\epsilon_{1}})}{n\,\epsilon_{1}(\epsilon+p\,\epsilon_{1})}-\frac{e^{-t^{1}\epsilon_{2}}(1-e^{-n\,t^{2}\epsilon_{2}})}{n\,\epsilon_{2}(\epsilon+p\,\epsilon_{2})}\right)\ ,$ (275)

### v4

@@ 7.1 Minkowski vacuum and the 4d prepotential
We briefly highlight key aspects of 4d ${\cal N}=2$ supergravity to clarify its connection to the equivariant topological strings framework. The 4d action, whether in superspace or component form, is governed by the prepotential, a holomorphic function of the complex scalars $X^{I}$.  [^34] In off-shell supergravity, the scalars $X^{I}$ are the lowest components of the vector multiplets and the vector superfield in superspace. On-shell, they overparametrize the physical scalars but are still used to define the theory via the prepotential. Supersymmetry ensures that the prepotential fully determines the Lagrangian, including higher-derivative corrections. It must also be a homogeneous function of a specific degree, related to the number of derivatives.
@@
where $C_{abc}$ are the triple intersection numbers of the CY three-fold, cf. (15) - (16), with $a,b,c=1,\dots,h^{1,1}(\mathrm{CY}_{3})$, and we assume an implicit summation over repeated up-down indices. The additional scalar $X^{0}$ (we use the convention $I=\{0,a\}$) in the denominator comes precisely from the circle reduction from 5d to 4d, and gives us a strong hint about the relation to equivariant topological strings on $\mathbb{C}\times\mathrm{CY}_{3}$. First, we can evaluate the equivariant volume of $X$, which clearly has a product structure
@@
We move on to the second term in the derivative expansion of the prepotential, which supersymmetry constrains to be of degree 0 in the scalars $X^{I}$ and linear in an additional composite scalar $\hat{A}$,  [^36] For more details on the superconformal formalism that suitably accounts for higher-derivative corrections, see e.g. Lauria:2020rhc . The scalar $\hat{A}$ in this case determines the coupling to the so-called Weyl2 four-derivative invariant, Bergshoeff:1980is .
$F^{4\partial}(X;\hat{A})=\frac{1}{24}\,\frac{b_{a}X^{a}}{X^{0}}\,\hat{A}\$, (235)
@@ 7.2 Black holes and the OSV conjecture
$\begin{array}[]{ccc}(\mathbb{C}\times\mathrm{CY}_{3})&\to&Y\\ &&\downarrow\\ &&\mathbb{P}^{1}\end{array}$ (241)
@@
Our equivariant approach proves useful here, as it explains why a formula like (242) should hold for the topological string from a 11d perspective. The following reformulation of the OSV conjecture highlights the unifying role of equivariant topological strings in connecting black hole entropy to geometric invariants.
@@
where we denoted the equivariant parameter on the two-sphere as $\bar{\nu}=a\epsilon_{-}-b\epsilon_{+}$ (with $a=b=1$ on the sphere), as we reserve the label 0 for the parameters coming from the $\mathbb{C}$ direction in the fiber as in the previous subsection. The equivariant volume of $\mathbb{C}\times\mathrm{CY}_{3}$ was already evaluated there, (232), and just as in Section 5 we denoted appropriate shift of the relevant parameters on the two poles with the $\pm$ subscript. On the complex plane we explicitly derived the shifts in $\lambda^{0}$ and $\epsilon_{0}$ in Subsection 5.1:
@@
The above conjecture for the shifts $t^{a}_{\pm}$ can be viewed as a reformulation of the OSV conjecture. It seems plausible that the shifts $t^{a}_{\pm}$ here come from the complexification of the real Kähler parameters by the reduction of the 11d three-form field, $C_{3}$. On the other hand, we have rigorously derived the shift along the toric direction $\epsilon_{0}^{\pm}$, finding perfect agreement with higher-derivative supergravity.
@@ 8 Discussion and outlook
We finish with a discussion that relates directly to the topological string partition function discussed here. We give a list of topics more closely related to supergravity and holography in Cassia:2025jkr .
@@
$F_{X}(\lambda,\epsilon;g_{s},\mathfrak{b})=\sum_{\mathfrak{g}=0}^{\infty}\sum_{n=0}^{\mathfrak{g}}(-1)^{n}g_{s}^{2(\mathfrak{g}-1)}\,(\mathfrak{b}+\mathfrak{b}^{-1})^{2n}F_{\mathfrak{g},n}(\lambda,\epsilon)\ ,$ (255)
@@
An apparent caveat of the refinement is the lack of a clear geometric origin for the additional parameter at the level of the CY manifold $X$. This particularly affects the constant maps terms, whose refinement lacks a general definition. We hope to come back to this point in the future, Cassia:2025jkr .
@@
where the power of $\omega$ gives the homogeneity degree in the moduli $\lambda^{i}$. Each of these three candidates correspond to one of the terms in (56). This is a choice that appears specifically relevant in the geometric description of M-theory, but this is no longer true for other brane models, see in particular (Colombo:2023fhu, , (1.7)).
@@
where $N_{\rm M2}$ and $N_{\rm M5}$ correspond to the exact brane charge of the corresponding type. We should however note that the relation between the transverse space $L$ and the modeled $X$ might generically become more complicated, see some examples in Martelli:2023oqk .
Similarly, one could imagine a different equivariant topological string definition, relating to double intersections numbers, where the constant maps terms will be related to the equivariant upgrade of
@@
which would be used in the description of D3-branes and other type IIB objects. Based on the scaling argument in (Colombo:2023fhu, , (1.7)), one would have to consider an integral of the form
@@
where the leading term in $F^{(2)}_{X}$ is proportional to $\mathbb{V}^{(2)}_{X}(\lambda,\epsilon)$, cf. (45), instead of $\mathbb{V}^{(3)}_{X}(\lambda,\epsilon)$, thus leading to a Gaussian integral in the variable $\lambda$. It then follows that the saddle-point equations for this integral match with the flux constraints considered in (Colombo:2023fhu, , (1.4)). At present, this should be viewed merely as a reformulation of the flux constraints in terms of a saddle-point equation for an integral over $\lambda$. It would nevertheless be interesting to understand whether there exists an a priori justification for such an integral representation, as in (259).
Likewise, we can consider an arbitrary degree $m$ in $\lambda$,
@@
with each higher degree allowing a larger number of different invariants made up from $m$ derivatives acting on the equivariant volume $\mathbb{V}_{X}$ (we skipped $c_{1}(TX)$ as it vanishes on CY manifolds non-equivariantly), and first term proportional to $\mathbb{V}_{X}^{(m)}$. Such an example of higher degree $({\rm deg}=5)$ is a candidate for the description of massive IIA objects, see again (Colombo:2023fhu, , (1.7)),
@@
Note furthermore that each of the terms of a given degree might allow for additional refinement and come in a different linear combination, see the discussion of periods in Cassia:2022lfj . We give additional holographic comments on these observations in Cassia:2025jkr .
@@ Acknowledgements
We wish to thank Thorsten Hertl for illuminating discussions, and Dario Martelli and Alberto Zaffaroni for useful comments on the draft. We gratefully acknowledge support during the MATRIX Program “New Deformations of Quantum Field and Gravity Theories,” (Creswick AU, 22 Jan – 2 Feb 2024), where this project was initiated. The work of L.C. was supported by the ARC Discovery Grant DP210103081. The study of K.H. is financed by the European Union-NextGenerationEU, through the National Recovery and Resilience Plan of the Republic of Bulgaria, project No BG-RRP-2.004-0008-C01.
Appendix A Exact M2-brane charge
Here we address the question of the exact M2-brane charge, denoted by $N_{\rm M2}$ in the introduction, first focusing on the spacetime filling case. In the classical two-derivative supergravity background, see section 6.1, $N_{\rm M2}$ matches the number of branes $N$. Eight-derivative corrections to the M2-brane solution were previously discussed in Bergman:2009zh , where two correction terms were proposed (in the absence of fractional branes). However, we find that only one of these terms should be included, while the second should be discarded.
To present our main argument, we briefly summarize the key elements of the calculation in Bergman:2009zh , which is based on the Chern-Simons-like higher-derivative correction to the 11d supergravity action, derived in Vafa:1995fj ; Duff:1995wd .  [^45] It is important to note that the fully supersymmetric invariant Lagrangian that includes this higher-derivative correction is still unknown. Hence, we only rely on the available action to determine the M2-brane charge, which is protected from other contributions, but cannot use it to calculate the corrections to the supergravity action. Including this correction, the 11d supergravity bosonic action for the metric $g_{\mu\nu}$ and the three-form field $C_{3}$ with field strength $F_{4}={\rm d}C_{3}$ is given by
$S_{\mathrm{11d}}=\frac{2\pi}{(2\pi l_{P})^{9}}\int\left[R\star 1-\frac{1}{2}\,F_{4}\wedge\star F_{4}-\frac{1}{6}\,C_{3}\wedge F_{4}\wedge F_{4}+(2\pi l_{P})^{6}\,C_{3}\wedge I_{8}\right]\ ,$ (262)
where $I_{8}$ is the 8-dimensional anomaly polynomial, which can be expressed as a polynomial in the Pontryagin classes of the tangent bundle to the 11d spacetime,
$I_{8}=-\frac{1}{2\cdot 4!}\,\left(\frac{1}{4}\,p_{1}^{2}-p_{2}\right)\ .$ (263)
The M2-brane charge in a given background can be derived from the equation of motion for $C_{3}$,
${\rm d}\star F_{4}+\frac{1}{2}\,F_{4}\wedge F_{4}=(2\pi l_{P})^{6}\,N\,\delta^{8}(x)+(2\pi l_{P})^{6}\,I_{8}\ ,$ (264)
where the first term on the r.h.s. represents a localized M2-brane source with flux $N$. This flux appears naturally in the explicit supergravity background solution discussed in Sec. 6, where the eight-dimensional space transverse to the M2-brane is identified with the CY cone $X_{\circ}$ over the Sasakian manifold $L$.
At large radius, the 11d geometry looks like that of $\mathbb{R}^{3}\times X_{\circ}$, so that the tangent bundle decomposes as the direct sum of a trivial rank-3 vector bundle and the tangent of $X_{\circ}$ (pulled-back to 11d). Since $X_{\circ}$ is a complex manifold with vanishing first Chern class (i.e. CY), we can further simplify the anomaly polynomial as
$I_{8}=-\frac{c_{4}(TX_{\circ})}{24}\ ,$ (265)
where $c_{4}$ is the fourth Chern class of the tangent to $X_{\circ}$, and the flat directions do not give any contributions.
The key point, which holds universally for all M2-brane solutions, is that the additional terms in the $C_{3}$ equation of motion generate corrections to the brane charge:
$N_{\rm M2}$= $\,\frac{1}{(2\pi l_{P})^{6}}\,\int_{L}\star F_{4}$ (266)
$=$ $\,\frac{1}{(2\pi l_{P})^{6}}\,\int_{X_{\circ}}\mathrm{d}\star F_{4}$
$=$ $\,N-\frac{\chi(X)}{24}-\frac{1}{2(2\pi l_{P})^{6}}\,\int_{X}F_{4}\wedge F_{4}\ ,$
where we used the equation of motion (264) together with Stokes’ theorem and identified $L$ with the boundary of $X_{\circ}$ at infinity in the radial direction. In the last line, we also used that the top Chern class of a smooth manifold integrates to the Euler characteristic. However, since $X_{\circ}$ is not smooth, we define the integral of the top Chern class by first resolving the conical singularity and then taking the mesonic limit $t\to 0$. We denote the resulting integral by $\chi(X)$ for short, and it is independent of the chosen resolution provided that the resolution contains no orbifold points.
The last term here accounts for a contribution from discrete torsion, arising from the discrete holonomy of the $C_{3}$ field. This term plays an important role in theories like ABJ Aharony:2008gk , where the ranks of the gauge groups in the quiver field theory do not match, see below.
Setting aside the discrete torsion term for the moment, we wish to highlight an important distinction in our interpretation of the correction proportional to $\chi(X)$, compared to that of Bergman:2009zh . In Bergman:2009zh , the authors argue that this topological term should be separated into two independent, non-vanishing contributions—a bulk and a boundary term. We find this distinction unnecessary, as the Euler characteristic is a topological invariant only when considered in its entirety, whereas the bulk and boundary terms are not separately metric independent. Since the M2-brane charge itself must be an invariant quantity, it is logical to include the full Euler characteristic on the r.h.s. of (266), as we have done.
This represents our only disagreement with the analysis in Bergman:2009zh , and we will use the M2-brane charge as presented here, in the absence of discrete torsion,
$N_{\rm M2}=N-\frac{\chi(X)}{24}\ .$ (267)
The contribution from discrete torsion—previously discussed in Bergman:2009zh in the context of the $\mathbb{C}^{4}/\mathbb{Z}_{k}$ orbifold—is instead associated with fractional brane charges Aharony:2008gk . In such cases, the gauge group ranks at the various nodes of the quiver theory become unequal, with the imbalance captured precisely by the discrete torsion term in (266). For the $\mathbb{C}^{4}/\mathbb{Z}_{k}$ example, our analysis is fully consistent with the results of Bergman:2009zh . A systematic generalization of the discrete torsion contribution to an arbitrary toric manifold $X$ remains an interesting open direction for future work.
Wrapped brane charges
In the case of wrapped M2-branes on a two-dimensional surface $\Sigma$, at two-derivative level, we have
$n_{i}\,N=\frac{1}{(2\pi l_{P})^{6}}\int_{S_{i}}\star F_{4}\in\mathbb{Z}\ ,$ (268)
where $S_{i}\hookrightarrow M$ are defined in (221). In the case that $\Sigma$ is a toric Riemann surface, there are two additional 7-dimensional submanifolds of $M$ that can support M2-brane charge; they correspond to the fibers of $M$ over the two fixed-points in the base. We denote these as $S_{\pm}$ and , by definition, they are isomorphic to two copies of the fiber $L$. Then we have two equations analogous to (268),
$aN=\frac{1}{(2\pi l_{P})^{6}}\int_{S_{+}}\star F_{4}\ ,\hskip 30.0pt-bN=\frac{1}{(2\pi l_{P})^{6}}\int_{S_{-}}\star F_{4}\ ,$ (269)
where we identified the base of the fibration with $\mathbb{WP}^{1}_{a,b}$. Due to the CY condition on the total space $Y$, we have that the sum of all toric divisors should vanish identically. i.e. $\sum_{i}S_{i}+S_{+}+S_{-}=0$, and in fact one can check that, because of (161), the sum of the integrals of $\star F_{4}$ over all divisors vanishes.
Following the previous discussion, we find a shift in the exact charges,
$N_{{\rm M2},i}:=n_{i}\,N+\delta_{i}\ ,\quad N_{{\rm M2},+}:=-a\,N+\delta_{+}\ ,\quad N_{{\rm M2},-}:=-b\,N+\delta_{-}\ ,$ (270)
given by
$\delta_{i}:=\int_{S_{i}}\star f_{4}\ ,\quad\delta_{\pm}:=\int_{S_{\pm}}\star f_{4}\ ,\qquad\mathrm{d}\star f_{4}=I_{8}\ ,$ (271)
again looking at the torsionless case for simplicity.
Let us assume there exist submanifolds $y_{i}$ of the 11d background geometry such that $\partial y_{i}=S_{i}$, then we can write $\delta_{i}=\int_{y_{i}}I_{8}$. A thorough analysis of the geometry of the $y_{i}$’s is beyond the scope of this article, however it natural to expect that $y_{i}$ take the form of the total spaces of fiber bundles over $\Sigma$ with fibers isomorphic to the toric divisors $D_{i}$ in $X_{\circ}$. Whenever the space $Y$ is itself a toric CY (and $\Sigma$ is a weighted projective space $\mathbb{WP}^{1}_{a,b}$), then it follows that $y_{i}$ are toric divisors in $Y$, corresponding to the vanishing of one of the homogeneous coordinates in the fiber. In addition to these divisors, $Y$ also has two divisors associated to the vanishing of the homogeneous coordinates on the base $\mathbb{WP}^{1}_{a,b}$, and they correspond to the fibers over the two fixed points in $\mathbb{WP}^{1}_{a,b}$; we denote these divisors as $y_{\pm}$ which correspond to the cones over $S_{\pm}$, respectively. In this case then, the shift in the brane charges can be written as
$\delta_{i}=-\frac{1}{24}\int_{y_{i}}c_{4}(TY)\,.$ (272)
Notice that the integrand is no longer the top Chern class of the space over which we integrate, which means that the result is not just the Euler characteristic of $y_{i}$ but something more complicated that depends on the topology of the fibration. Similarly, for $y_{\pm}$ we have
$\delta_{\pm}=-\frac{1}{24}\int_{y_{\pm}}c_{4}(TY)\,.$ (273)
Appendix B $\mathbb{P}^{k}$ and $\mathbb{C}^{m}\times\mathbb{P}^{k}$ fibrations
@@
$Q=\begin{pmatrix}1&1&0&-k\\ 0&0&1&1\end{pmatrix}\$, (274)
@@
$Q=\begin{pmatrix}a&b&0&-k\\ 0&0&1&1\end{pmatrix}\$, (282)
@@
$Q=\begin{pmatrix}a&b&0&\dots&0&-k\\ 0&0&1&\dots&1&1\end{pmatrix}\$, (284)
@@
$\mathbb{C}\times\mathbb{P}^{1}$ over $\mathbb{P}^{1}$
Let us consider the fiber $\mathbb{C}\times\mathbb{P}^{1}$ over a $\mathbb{P}^{1}$ base, giving us an interesting example of a product space fiber with a compact and a non-compact factor. This is the closest example to the situation discussed in the main text in section 7.2 that features a non-toric compact factor in the fiber.
@@
$Q=\begin{pmatrix}1&1&0&-n&-p\\ 0&0&1&1&0\end{pmatrix}\$, (286)
@@
It is particularly interesting to observe what happens in the non-equivariant limits in this case. Setting to zero the equivariant parameters $\epsilon_{3},\epsilon_{4}$ associated to the compact directions on the fiber, we find
$\mathrm{vol}(t,\epsilon,\epsilon_{1,2},\epsilon_{3,4}=0)=\frac{1}{\epsilon_{2}-\epsilon_{1}}\,\left(\frac{\mathrm{e}^{-t^{1}\epsilon_{1}}(1-\mathrm{e}^{-n\,t^{2}\epsilon_{1}})}{n\,\epsilon_{1}(\epsilon+p\,\epsilon_{1})}-\frac{\mathrm{e}^{-t^{1}\epsilon_{2}}(1-\mathrm{e}^{-n\,t^{2}\epsilon_{2}})}{n\,\epsilon_{2}(\epsilon+p\,\epsilon_{2})}\right)\ ,$ (289)